# Muon Optimizes Under Spectral Norm Constraints

### v1

@@ 7.4 Convergence of Muon to a KKT point
Let $\{\mathbf{X}_{t}\}_{t\in\mathbb{N}}$ be a stochastic process. A set $\mathcal{M}$ is called the $\omega$-limit set of the process if it is the minimal closed set satisfying
$\Pr\left(\lim_{t\to\infty}d(\mathbf{X}_{t},\mathcal{M})=0\right)=1.$
In other words, $\mathcal{M}$ is the smallest closed set such that the trajectories approach $\mathcal{M}$ a.s. as $t\to\infty$. This is equivalent to $\mathcal{M}$ being the support of the union of all limit measures of $\{\mathbf{X}_{t}\}_{t\in\mathbb{N}}$.
@@
Lemma 9 (LaSalle’s invariance principle for stochastic dynamical systems [ZLW22]).
Let $\{\mathbf{X}_{t}\}_{t\in\mathbb{N}}$ be a stochastic process contained in a bounded set a.s., and suppose there exist a nonnegative function $\mathcal{V}$ and a nonnegative, lower semicontinuous continuous function $h$ such that
[…]
@@
Let $\mathcal{M}$ be the $\omega$-limit set of $\{\mathbf{X}_{t}\}_{t\in\mathbb{N}}$. Then $\mathcal{M}$ is contained in the set $\{\mathbf{X}\in\mathbb{X}\mid h(\mathbf{X})=0\}$.
@@
Under Assumptions 1, 2, and 3, let $10<\beta_{1}<\beta_{2}<$ 1, $\lambda>0$, $\eta_{t}<\frac{1}{\lambda}$, and $\sum_{t=0}^{\infty}\eta_{t}=\infty$, and suppose $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (2). Then $\mathbf{X}_{t}$ converges to
[…]
@@
[…]
by Proposition 2. We conclude that $\mathbf{X}_{t}$ converges to $\{\mathbf{X}\in\mathbb{X}\mid h(\mathbf{X})=0\}=\mathbb{B}$ a.s. by Lemma 9. ∎
@@
In the setting of Theorem 4, let $\left\lVert\lambda\mathbf{X}_{0}\right\rVert_{\textup{op}}\leq 1$ and $\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty$. Then as $\frac{\sigma}{\sqrt{n_{\mathrm{batch}}}}\to 0$, $\mathbf{X}_{t}$ converges to a KKT point of (5) a.s.
@@
Since $\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty$, we have $\lim_{t\to\infty}\eta_{t}=0$. Since $\operatorname{msgn}(\widetilde{\mathbf{M}}_{t+1})-\lambda\mathbf{X}_{t+1}$ has bounded norm, this implies $\mathbf{X}_{t+1}-\mathbf{X}_{t}=\eta_{t}(\operatorname{msgn}(\widetilde{\mathbf{M}}_{t+1})-\lambda\mathbf{X}_{t+1})\to\mathbf{0}$, so $\mathbf{X}_{t}$ converges to a point $\mathbf{X}^{\star}$. Then by the smoothness of $\mathcal{F}$, we have that $\nabla\mathcal{F}(\mathbf{X}_{t})$ converges to $\nabla\mathcal{F}(\mathbf{X}^{\star})$, so
$(\mathbf{X}_{t},\mathbf{M}_{t},\widetilde{\mathbf{M}}_{t},\eta_{t})\to(\mathbf{X}^{\star},-\nabla\mathcal{F}(\mathbf{X}^{\star}),-\nabla\mathcal{F}(\mathbf{X}^{\star}),0).$ (26)
@@
$h(\mathbf{X},\mathbf{M},\widetilde{\mathbf{M}},\eta)$ $=\left\lVert\widetilde{\mathbf{M}}\right\rVert_{\textup{tr}}-\left\langle\lambda\mathbf{X},\widetilde{\mathbf{M}}\right\rangle,\quad\alpha_{t}=\eta_{t},\quad\text{and}\quad\gamma_{t}=2\eta_{t}^{2}C_{\mathcal{K}}^{2}L,$
@@
[…]
by (23) and $a_{t}\Gamma_{t}+b_{t}\Delta_{t}\geq\Gamma_{t}=h(\mathbf{Z}_{t+1})$. We conclude that $h(\mathbf{Z}^{\star})=0$ a.s., where $\mathbf{Z}^{\star}:=\lim_{t\to\infty}\mathbf{Z}_{t}$, by Lemma 9. Then (26) implies
[…]
so $\mathbf{X}^{\star}$ is a KKT point of (5) a.s. by Proposition 1. ∎
We remark that the feasible region $\mathbb{B}$ has nonempty interior, so Slater’s condition holds. This implies that the KKT conditions are necessary for optimality, and under the additional assumption that $\mathcal{F}$ is convex, the KKT conditions also become sufficient for optimality. In this case, Theorem 5 shows that Muon converges to an optimal solution of (5).
@@ 8.1 Toy example
Figure 2: Trajectories of the Muon optimizer for the matrix-valued optimization problem $\min_{\mathbf{X}\in\mathbb{R}^{2\times 2}}f(\mathbf{X}),\text{ where }f(\mathbf{X})=\left\lVert\mathbf{A}\mathbf{X}-\mathbf{B}\right\rVert_{\textup{F}}^{2}+\mu\left\lVert\mathbf{X}\right\rVert_{\textup{F}}^{2}$, $\mathbf{A},\mathbf{B}\in\mathbb{R}^{2\times 2}$, and $\mu\in\mathbb{R}_{>0}$, evaluated for two different values of $\lambda$: $1.5$ (blue) and $5.0$ (green). The colored boxes illustrate the constraint sets induced by $\|\mathbf{X}\|_{\infty}\leq\frac{1}{\lambda}$: the blue box corresponds to $\lambda=1.5$, and the green box corresponds to $\lambda=5.0$. The left plot shows the trajectories of the singular values across iterations, starting from two initial points $(1,0)$ and $(0,1)$ (black dots), which are equivalent due to symmetry. The red dot indicates the optimal solution.
@@ 8.3 Implicit spectral regularization in large models
We investigate the implicit spectral regularization induced by the Muon optimizer in comparison to AdamW. Figure 6 displays the singular value distributions of converged weights for the query ($\mathbf{W}_{\mathrm{Q}}$), key ($\mathbf{W}_{\mathrm{K}}$), and value ($\mathbf{W}_{\mathrm{V}}$) matrices in the LLaMA 0.5B model trained by both optimizers. We observe that Muon consistently produces regularized singular value distributions, reflecting the effect of its implicit nuclear norm constraint. In contrast, singular values from AdamW do not exhibit any spectral regularization.
@@ 8.4 Generalizations via different convex functions
Figure 7: Behavior of the Muon optimizer under different choices of $\mathcal{K}$. Using the same objective function as in Figure 2, varying the choice of $\mathcal{K}$ induces distinct implicit constraints and penalty structures on the objective function. The colored boxes illustrate the constraint sets induced by $\mathcal{K}$. For example, choosing the thresholding function $\mathcal{K}(\mathbf{X})=\sum_{i=1}^{\min(n,m)}\max(|\boldsymbol{\sigma}_{i}(\mathbf{X})|-e,0)$ enforces a soft-thresholding penalty on singular values $\boldsymbol{\sigma}_{i}(\mathbf{X})$ exceeding a threshold $e$. Here, $\left\lVert\boldsymbol{\sigma}_{i}(\mathbf{X})\right\rVert_{p}$ denotes the Schatten $p$-norm of $\mathbf{X}$.
@@ 9 Conclusion
In this paper, we showed that Muon is an instance of the Lion-$\mathcal{K}$ optimizer when equipped with the nuclear norm and extended the Lion-$\mathcal{K}$ analysis of [CLLL24] to accommodate matrix-valued updates in both deterministic and stochastic gradient settings. We tailored our analysis to Muon with decoupled weight decay, demonstrating that it converges to a KKT point of a spectral-norm-constrained optimization problem. Beyond theoretical guarantees, we developed a framework for generalizing Muon and empirically validated our results. Overall, we present Muon as a theoretically grounded optimizer for deep learning, with promising directions for future work.
@@
While our theoretical and empirical findings provide substantial insights, several limitations suggest directions for future research. First, although our analysis focuses on key convex maps such as the nuclear norm and clipping functions, extending the Lion-$\mathcal{K}$ framework to a broader class of convex maps may reveal additional implicit regularization behaviors tailored to specific tasks. Second, extending our results to practical training conditions (e.g. general learning rates, nonsmooth objectives) warrants further investigation. Finally, scaling empirical evaluations to larger models and more diverse tasks would help further validate and refine the practical applicability and robustness of the Muon optimizer and its Lion-$\mathcal{K}$ generalizations.
@@ Appendix A Continuous-time analysis of Lion-$\mathcal{K}$
Under Assumptions 1 and 4, let $\mathcal{F}$, $\mathcal{K}$, and $\mathcal{K}^{*}$ be continuously differentiable and
@@
00 0 $\geq\left\langle-\widetilde{\mathbf{M}}+\nabla\mathcal{K}^{*}(\lambda\mathbf{X}),\nabla\mathcal{K}(\widetilde{\mathbf{M}})-\lambda\mathbf{X}\right\rangle$ (29)
@@
00 0 $\geq\left\langle\mathbf{M}-\widetilde{\mathbf{M}},\nabla\mathcal{K}(\widetilde{\mathbf{M}})-\nabla\mathcal{K}(\mathbf{M})\right\rangle$
@@
$\phantom{=}+\frac{1-\epsilon\gamma}{1+\epsilon\lambda}\left\langle\nabla\mathcal{K}(\mathbf{M})-\lambda\mathbf{X},-\alpha\nabla\mathcal{F}(\mathbf{X})-\gamma\mathbf{M}\right\rangle$
@@
$\phantom{=}+\frac{1-\epsilon\gamma}{\epsilon(1+\epsilon\lambda)}\left\langle\epsilon\alpha\nabla\mathcal{F}(\mathbf{X})+\epsilon\gamma\mathbf{M},(\nabla\mathcal{K}(\widetilde{\mathbf{M}})-\lambda\mathbf{X})-(\nabla\mathcal{K}(\mathbf{M})-\lambda\mathbf{X})\right\rangle$

### v2

@@ 7.4 Convergence of Muon to the set of KKT points
Let $\{\mathbf{X}_{t}\}_{t\in\mathbb{N}}$ be a stochastic process. A set $\mathbb{M}$ is called the $\omega$-limit set of the process if it is the minimal closed set satisfying
$\Pr\left(\lim_{t\to\infty}d(\mathbf{X}_{t},\mathbb{M})=0\right)=1.$
In other words, $\mathbb{M}$ is the smallest closed set such that the trajectories approach $\mathbb{M}$ a.s. as $t\to\infty$. This is equivalent to $\mathbb{M}$ being the support of the union of all limit measures of $\{\mathbf{X}_{t}\}_{t\in\mathbb{N}}$.
@@
Lemma 9 (LaSalle’s invariance principle for stochastic dynamical systems).
Let $\{\mathbf{X}_{t}\}_{t\in\mathbb{N}}$ be a stochastic process contained in a bounded set a.s., and suppose there exist a nonnegative function $\mathcal{V}$ and a nonnegative, lower semicontinuous function $h$ such that
$\mathbb{E}[\mathcal{V}(\mathbf{X}_{t+1})\mid\mathcal{F}_{t}]-\mathcal{V}(\mathbf{X}_{t})\leq-\alpha_{t}h(\mathbf{X}_{t+\ell})+\gamma_{t}\text{ a.s.},$
@@
Let $\mathbb{M}$ be the $\omega$-limit set of $\{\mathbf{X}_{t}\}_{t\in\mathbb{N}}$. Then $\mathbb{M}$ is contained in the set $\{\mathbf{X}\in\mathbb{X}\mid h(\mathbf{X})=0\}$.
@@
Under Assumptions 1, 2, and 4, let $0\leq\beta_{1}<\beta_{2}<1$, $\lambda>0$, $\eta_{t}\leq\frac{1}{\lambda}$, and $\sum_{t=0}^{\infty}\eta_{t}=\infty$, and suppose $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (2). Then $\mathbf{X}_{t}$ converges to
$\mathbb{B}:=\{\mathbf{X}\in\mathbb{X}\mid\left\lVert\lambda\mathbf{X}\right\rVert_{\textup{op}}\leq 1\}\text{ a.s.}$
@@
$\mathbb{E}[\mathcal{V}(\mathbf{X}_{t+1})\mid\mathbf{X}_{t}]-\mathcal{V}(\mathbf{X}_{t})=\mathbb{E}\left[\left\lVert\mathbf{X}_{t+1}\right\rVert_{\textup{op}}-\frac{1}{\lambda}\;\middle|\;\mathbf{X}_{t}\right]-\left(\left\lVert\mathbf{X}_{t}\right\rVert_{\textup{op}}-\frac{1}{\lambda}\right)\leq-\alpha_{t}h(\mathbf{X}_{t})+\gamma_{t}\text{ a.s.}$
by Proposition 1. Thus $\mathbf{X}_{t}$ converges to $\{\mathbf{X}\in\mathbb{X}\mid h(\mathbf{X})=0\}=\mathbb{B}$ a.s. by Lemma 9. ∎
@@
In the setting of Theorem 5, let $\left\lVert\lambda\mathbf{X}_{0}\right\rVert_{\textup{op}}\leq 1$, $\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty$, and
$\sum_{t=0}^{\infty}\eta_{t}\left[\frac{\sigma^{2}}{n_{\mathrm{batch}}}\right]_{t}^{\frac{1}{2}}<\infty.$
Then $\mathbf{X}_{t}$ converges to the set of KKT points of (5) a.s.
@@
$h(\mathbf{X},\mathbf{M},\widetilde{\mathbf{M}},\eta)$ $=\left\lVert\widetilde{\mathbf{M}}\right\rVert_{\textup{tr}}-\left\langle\lambda\mathbf{X},\widetilde{\mathbf{M}}\right\rangle,\quad\alpha_{t}=\eta_{t},\quad\text{and}\quad\gamma_{t}=2\eta_{t}^{2}C_{\mathcal{K}}^{2}L+\eta_{t}C_{\mathcal{K}}\left[\frac{\sigma^{2}}{n_{\mathrm{batch}}}\right]_{t}^{\frac{1}{2}},$
@@
$\mathbb{E}[\mathcal{H}(\mathbf{Z}_{t+1})\mid\mathbf{Z}_{t}]-\mathcal{H}(\mathbf{Z}_{t})\leq-\alpha_{t}(a_{t}\Gamma_{t}+b_{t}\Delta_{t})+\gamma_{t}\leq-\alpha_{t}h(\mathbf{Z}_{t+1})+\gamma_{t}\text{ a.s.}$
by (24) and $a_{t}\Gamma_{t}+b_{t}\Delta_{t}\geq\Gamma_{t}=h(\mathbf{Z}_{t+1})$. Thus by Lemma 9, $h(\mathbf{Z}_{t})=0$ a.s. as $t\to\infty$, i.e.
$\lim_{t\to\infty}\left\lVert\widetilde{\mathbf{M}}_{t}\right\rVert_{\textup{tr}}-\left\langle\lambda\mathbf{X}_{t},\widetilde{\mathbf{M}}_{t}\right\rangle=0\text{ a.s.}$
It follows that, with probability 1,
$\lim_{t\to\infty}\mathcal{S}(\mathbf{X}_{t})$ $=\lim_{t\to\infty}\left(\left\lVert\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}-\widetilde{\mathbf{M}}_{t}\right\rVert_{\textup{tr}}+\left\langle\lambda\mathbf{X}_{t},\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}-\widetilde{\mathbf{M}}_{t}\right\rangle\right)$
$\leq\lim_{t\to\infty}\left(\left\lVert\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}\right\rVert_{\textup{tr}}+\left\lVert\widetilde{\mathbf{M}}_{t}\right\rVert_{\textup{tr}}+\left\langle\lambda\mathbf{X}_{t},\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}\right\rangle-\left\langle\lambda\mathbf{X}_{t},\widetilde{\mathbf{M}}_{t}\right\rangle\right)$
$\leq\lim_{t\to\infty}\left(\left\lVert\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}\right\rVert_{\textup{tr}}+\left\lVert\lambda\mathbf{X}_{t}\right\rVert_{\textup{op}}\left\lVert\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}\right\rVert_{\textup{tr}}\right)\leq\lim_{t\to\infty}2\left\lVert\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}\right\rVert_{\textup{tr}}$
$\leq\lim_{t\to\infty}\Bigg(4\eta_{t-1}C_{\mathcal{K}}^{2}L+4\beta_{1}C_{\mathcal{K}}^{2}L\sum_{k=1}^{t-1}\beta_{2}^{t-k-1}\eta_{k-1}+2\beta_{1}\beta_{2}^{t-1}C_{\mathcal{K}}\left\lVert\nabla\mathcal{F}(\mathbf{X}_{0})+\mathbf{M}_{0}\right\rVert_{\textup{F}}$
$\quad+2(1-\beta_{1})\left\lVert\mathbf{G}_{t-1}-\nabla\mathcal{F}(\mathbf{X}_{t-1})\right\rVert_{\textup{tr}}+2\beta_{1}(1-\beta_{2})\sum_{k=0}^{t-2}\beta_{2}^{t-k-2}\left\lVert\mathbf{G}_{k}-\nabla\mathcal{F}(\mathbf{X}_{k})\right\rVert_{\textup{tr}}\Bigg)$
$=0,$
where the fourth line uses (31), the fifth line vanishes because of the asymptotically deterministic gradient, and the sixth line uses the Silverman–Toeplitz theorem to show that the sums vanish. By Proposition 2, we conclude that $\mathbf{X}_{t}$ converges to the set of KKT points of (5) a.s. ∎
Remark 1.
The feasible region $\mathbb{B}$ has nonempty interior, so Slater’s condition holds. This implies that the KKT conditions are necessary for optimality, and under the additional assumption that $\mathcal{F}$ is convex, the KKT conditions also become sufficient for optimality. In this case, Theorem 6 shows that the algorithm converges to the set of optimal solutions of (5).
Remark 2.
The $\sum_{t=0}^{\infty}\eta_{t}\left[\frac{\sigma^{2}}{n_{\mathrm{batch}}}\right]_{t}^{\frac{1}{2}}<\infty$ condition in Theorem 6 can be achieved by using deterministic gradients or sufficiently increasing batch sizes, e.g. $n_{\mathrm{batch}}]_{t}=\Theta(t)$ when $\eta_{t}=O(t^{-1})$.
Remark 3.
Theorem 6 does not guarantee convergence to a single point; compare with Frank–Wolfe, which can fail to converge even in the smooth convex setting [BCP24].
@@ 8.1 Toy example
Figure 2: Trajectories of Muon for the matrix optimization problem $\min_{\mathbf{X}\in\mathbb{R}^{2\times 2}}f(\mathbf{X})$, where $f(\mathbf{X})=\left\lVert\mathbf{A}\mathbf{X}-\mathbf{B}\right\rVert_{\textup{F}}^{2}+\mu\left\lVert\mathbf{X}\right\rVert_{\textup{F}}^{2}$, $\mathbf{A},\mathbf{B}\in\mathbb{R}^{2\times 2}$, and $\mu\in\mathbb{R}_{>0}$, evaluated for two different values of $\lambda$: $1.5$ (blue) and $5.0$ (green). The colored boxes illustrate the constraint sets induced by $\|\mathbf{X}\|_{\infty}\leq\frac{1}{\lambda}$: the blue box corresponds to $\lambda=1.5$, and the green box corresponds to $\lambda=5.0$. The red dot indicates the optimal solution.
@@ 8.3 Implicit spectral regularization in large models
We investigate the implicit spectral regularization induced by the Muon optimizer in comparison to AdamW. Figure 6 displays the singular value distributions of converged weights for the query ($\mathbf{W}_{\mathrm{Q}}$), key ($\mathbf{W}_{\mathrm{K}}$), and value ($\mathbf{W}_{\mathrm{V}}$) matrices in the LLaMA 0.5B model trained by both optimizers. We observe that Muon consistently produces regularized singular value distributions, reflecting the effect of its implicit spectral norm constraint. In contrast, singular values from AdamW do not exhibit any spectral regularization.
@@ 8.4 Generalizations via different convex functions
Figure 7: Behavior of the Lion-$\mathcal{K}$ optimizer under different choices of $\mathcal{K}$. Using the same objective function as in Figure 2, varying the choice of $\mathcal{K}$ induces distinct implicit constraints and penalty structures on the objective function. The colored boxes illustrate the constraint sets induced by $\mathcal{K}$. For example, choosing the thresholding function $\mathcal{K}(\mathbf{X})=\sum_{i=1}^{\min(n,m)}\max(|\boldsymbol{\sigma}_{i}(\mathbf{X})|-e,0)$ enforces a soft-thresholding penalty on singular values $\boldsymbol{\sigma}_{i}(\mathbf{X})$ exceeding a threshold $e$. Here, $\left\lVert\boldsymbol{\sigma}_{i}(\mathbf{X})\right\rVert_{p}$ denotes the Schatten $p$-norm of $\mathbf{X}$.
@@ 9 Conclusion
In this paper, we showed that Muon is an instance of the Lion-$\mathcal{K}$ optimizer when equipped with the nuclear norm and extended the Lion-$\mathcal{K}$ analysis of [CLLL24] to accommodate matrix-valued updates in both deterministic and stochastic gradient settings. We tailored our analysis to Muon with decoupled weight decay, demonstrating that it converges to the set of KKT points of a spectral-norm-constrained optimization problem. Beyond theoretical guarantees, we developed a framework for the generalization of Muon and empirically validated our results. Overall, we present Muon as a theoretically grounded optimizer for deep learning, with promising directions for future work.
@@
While our theoretical and empirical findings provide substantial insights, several limitations suggest directions for future research. First, although our analysis focuses on key convex maps, such as the nuclear norm and clipping functions, extending the Lion-$\mathcal{K}$ framework to a broader class of convex maps may reveal additional implicit regularization behaviors tailored to specific tasks. Second, extending our results to practical training conditions (e.g. general learning rates, nonsmooth objectives) warrants further investigation. Finally, scaling empirical evaluations to larger models and more diverse tasks would help further validate and refine the practical applicability and robustness of the Muon optimizer and its Lion-$\mathcal{K}$ generalizations.
@@ Appendix A Continuous-time analysis of Lion-$\mathcal{K}$
Under Assumptions 1 and 5, let $\mathcal{F}$, $\mathcal{K}$, and $\mathcal{K}^{*}$ be continuously differentiable and
@@
$0$ $\geq\left\langle-\widetilde{\mathbf{M}}+\nabla\mathcal{K}^{*}(\lambda\mathbf{X}),\nabla\mathcal{K}(\widetilde{\mathbf{M}})-\lambda\mathbf{X}\right\rangle$ (28)
@@
$0$ $\geq\left\langle\mathbf{M}-\widetilde{\mathbf{M}},\nabla\mathcal{K}(\widetilde{\mathbf{M}})-\nabla\mathcal{K}(\mathbf{M})\right\rangle$
@@
$\quad+\frac{1-\epsilon\gamma}{1+\epsilon\lambda}\left\langle\nabla\mathcal{K}(\mathbf{M})-\lambda\mathbf{X},-\alpha\nabla\mathcal{F}(\mathbf{X})-\gamma\mathbf{M}\right\rangle$
@@
$\quad+\frac{1-\epsilon\gamma}{\epsilon(1+\epsilon\lambda)}\left\langle\epsilon\alpha\nabla\mathcal{F}(\mathbf{X})+\epsilon\gamma\mathbf{M},(\nabla\mathcal{K}(\widetilde{\mathbf{M}})-\lambda\mathbf{X})-(\nabla\mathcal{K}(\mathbf{M})-\lambda\mathbf{X})\right\rangle$